### v1

# Recurrent Interpolants for Probabilistic Time Series Prediction

Abstract
Sequential models such as recurrent neural networks or transformer-based models became de facto tools for multivariate time series forecasting in a probabilistic fashion, with applications to a wide range of datasets, such as finance, biology, medicine, etc. Despite their adeptness in capturing dependencies, assessing prediction uncertainty, and efficiency in training, challenges emerge in modeling high-dimensional complex distributions and cross-feature dependencies. To tackle these issues, recent works delve into generative modeling by employing diffusion or flow-based models. Notably, the integration of stochastic differential equations or probability flow successfully extends these methods to probabilistic time series imputation and forecasting. However, scalability issues necessitate a computational-friendly framework for large-scale generative model-based predictions. This work proposes a novel approach by blending the computational efficiency of recurrent neural networks with the high-quality probabilistic modeling of the diffusion model, which addresses challenges and advances generative models’ application in time series forecasting. Our method relies on the foundation of stochastic interpolants and the extension to a broader conditional generation framework with additional control features, offering insights for future developments in this dynamic field.
Preprint Equal Contribution Work done as part of an internship at Morgan Stanley
1 Introduction
Autoregression models [Box et al., 2015], such as recurrent neural networks [Graves, 2013, Sutskever et al., 2014, Hochreiter and Schmidhuber, 1997] or transformer models [Vaswani et al., 2017], have been the go-to methods for time series forecasting. They are widely applied in finance, biological statistics, medicine, geophysical applications, etc., effectively showcasing their ability to capture short-term and long-term sequential dependencies [Morrill et al., 2021]. These methods can also provide an assessment of prediction uncertainty through probabilistic forecasting by incorporating specific parametric probabilistic models into the output layer of the neural network. For instance, a predictor can model the Gaussian distribution by predicting both mean and covariance. However, the probabilistic output layer is confined within a simple probability family because the density needs to be parameterized by neural networks, and the loss must be differentiable with respect to neural network parameters.
To better capture sophisticated distributions in time series modeling and learn both the temporal and cross-feature dependencies, a common strategy involves exploring the generative modeling of time series using efficient distribution transportation plans, especially via diffusion or flow based models. For example, recent works such as Li et al. [2020] propose using latent neural SDE as latent state for modeling time series in a stochastic manner, while Spantini et al. [2022] summarize non-linear extensions of state space models using both deterministic and stochastic transformation plans. Tashiro et al. [2021], Biloš et al. [2023], Chen et al. [2023a], Miguel et al. [2022], Li et al. [2022] studied the application of diffusion models in probabilistic time series imputation and forecasting. The generative model is trained to learn the joint density of the whole time series window $\mathbf{x}\in\mathbb{R}^{K\times(T_{\mathrm{context}}+T_{\mathrm{prediction}})}$ with $K$ features. $T_{\mathrm{context}}$ is the size of the context window, $T_{\mathrm{prediction}}$ is the size of the prediction window. During inference, the model performs conditional generation given the context, similar to the inpainting task in computer vision [Song et al., 2021]. Compared to a recurrent model, where the model size is only proportional to the number of features, but not the length of the time window, such generative model predictors may suffer from scalability issues because the model size is related to both feature dimension and the size of the window. A more computational friendly framework is needed for large scale generative model-based time series prediction problems.
Generative modeling excels at modeling complicated high dimension distributions, but most models require learning a mapping from noise distribution to data distribution. If the generative procedure starts from an initial distribution proximate to the terminal data distribution, it can remarkably alleviate learning challenges, reduce inference complexity, and enhance the quality of generated samples, which is also supported by previous studies [Rubanova et al., 2019, Rasul et al., 2021a, b, Chen et al., 2023b, Deng et al., 2024a, b]. Time series data is typically continuous and neighboring time points exhibit strong correlations, indicating that the distribution of future time points is close to that of the current time point.
These observations inspire the creation of a time series prediction model under the generative framework that maps between dependent data points: Initiating the prediction of future time point’s distribution with the current time point is more straightforward and yields better quality; meanwhile, the longer temporal dependency is encoded by a recurrent neural network and the embedded history is passed to the generative model as the guidance of the prediction for the future time points. The new framework benefits from the efficient training and computation inherited from the recurrent neural network, while enjoying the high quality of probabilistic modeling empowered by the diffusion model.
Our contributions include:
  • We first extend the theory of stochastic interpolants to more general conditional generation framework with extra control features.
  • We adopt a conditional stochastic interpolants module for the sequential modeling and time series prediction, which is computational-friendly and achieves high quality modeling of the future time point’s distribution.
2 Background
As we formalize probabilistic time series forecasting within the generative framework in Section 4, this section is dedicated to reviewing commonly used generative methods and their extensions for conditional generation. These models will serve as baseline models in subsequent sections. For a broader overview of time series forecasting problems, refer to [Salinas et al., 2019, Alexandrov et al., 2019], and the references therein.
2.1 Denoising Diffusion Probabilistic Model (DDPM)
DDPM [Sohl-Dickstein et al., 2015, Ho et al., 2020] adds Gaussian noise to the observed data point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ at different scales $0<\beta_{1}<\beta_{2}<\dots<\beta_{N}$ such that the first noisy value $\mathbf{x}_{1}$ is close to the clean data $\mathbf{x}_{0}$, and the final value $\mathbf{x}_{N}$ is indistinguishable from noise. The generative model learns to revert this process allowing sampling new points from noise.
Following previous convention, we define $\bar{\alpha}_{n}=\prod_{k=1}^{n}\alpha_{k}$, with $\alpha_{n}=1-\beta_{n}$. Then the transition kernel is Gaussian and can be computed directly from $\mathbf{x}_{0}$:
$q(\mathbf{x}_{n}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{n}}\mathbf{x}_{0},(1-\bar{\alpha}_{n}){\bm{I}}).$ (1)
The posterior distribution is available in closed form:
$q(\mathbf{x}_{n-1}|\mathbf{x}_{n},\mathbf{x}_{0})=\mathcal{N}(\tilde{{\bm{\mu}}}_{n},\tilde{\beta}_{n}{\bm{I}}),$ (2)
where $\tilde{{\bm{\mu}}}_{n}$ depends on $\mathbf{x}_{0}$, $\mathbf{x}_{n}$ and a choice of $\beta$-scheduler. The generative model $p(\mathbf{x}_{n-1}|\mathbf{x}_{n})\approx q(\mathbf{x}_{n-1}|\mathbf{x}_{n},\mathbf{x}_{0})$ approximates the reverse process. The actual model $\epsilon_{\theta}(\mathbf{x}_{0},n)$ is usually reparameterized to predict the noise added to a clean data point, from the noisy data point $\mathbf{x}_{n}$. The loss function can be simply written as:
$\mathcal{L}=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}({\bm{0}},{\bm{I}}),n\sim\mathcal{U}(\{1,\dots,N\})}\left[\lVert\epsilon_{\theta}(\mathbf{x}_{n},n)-\bm{\epsilon}\rVert_{2}^{2}\right].$ (3)
Sampling new data is performed by first sampling a point from the pure noise $\mathbf{x}_{N}\sim\mathcal{N}({\bm{0}},{\bm{I}})$ and then gradually denoising it using the above model to get a sample from the data distribution [Ho et al., 2020].
2.2 Score-based Generative Model (SGM)
SGM [Song et al., 2021], like DDPM, considers a pair of forward and backward dynamics between $s\in[0,1]$:
${\mathrm{d}}\mathbf{x}_{s}=$ $f(\mathbf{x}_{s},s){\mathrm{d}}s+g(s){\mathrm{d}}\textbf{w}_{s}$ (4)
${\mathrm{d}}\mathbf{x}_{s}=$ $[f(\mathbf{x}_{s},s)-g(s)^{2}\nabla_{\mathbf{x}_{s}}\log p(\mathbf{x}_{s})]{\mathrm{d}}s+g(s){\mathrm{d}}\textbf{w}_{s}$ (5)
where $\nabla_{\mathbf{x}_{s}}\log p(\mathbf{x}_{s})$ is the so-called score function. The forward process usually is scheduled as simple processes, such as Brownian motion or Ornstein–Uhlenbeck process, which can transport data distribution to standard Gaussian distribution. The generative process is achieved by the backward process that walks from Gaussian prior distribution to the data distribution of interest. Now, Equation 5 gives a way to generate new points by starting at $\mathbf{x}_{s}\sim\mathcal{N}({\bm{0}},{\bm{I}})$ and solving the SDE backward in time giving $\mathbf{x}_{0}$ as a sample from data distribution. In practice, the only missing piece is obtaining the score. A standard approach is to approximate the score with a neural network.
Since during training, we have access to clean data, the score function is available in closed form. The model $\epsilon_{\theta}(\mathbf{x}_{s},s)$ learns to approximate the score from noisy data only, resulting in a loss function similar to Equation 3:
$\begin{split}\mathcal{L}=&\mathbb{E}_{s\sim\mathcal{U}(0,1),\mathbf{x}_{0}\sim\text{Data},\mathbf{x}_{s}\sim p(\mathbf{x}_{s}|\mathbf{x}_{0})}\Big{[}\\ &\lVert\epsilon_{\theta}(\mathbf{x}_{s},s)-\nabla_{\mathbf{x}_{s}}\log p(\mathbf{x}_{s}|\mathbf{x}_{0})\rVert_{2}^{2}\Big{]}.\end{split}$ (6)
2.3 Flow Matching (FM)
Flow matching [Lipman et al., 2022] constructs a probability path by learning the vector field that generates it. Given a data point $\mathbf{x}_{1}$, the conditional probability path is denoted with $p_{s}(\mathbf{x}|\mathbf{x}_{1})$ for $s\in[0,1]$. We put the constraints on $p_{s}(\mathbf{x}|\mathbf{x}_{1})$ such that $p_{0}(\mathbf{x}|\mathbf{x}_{1})=\mathcal{N}({\bm{0}},{\bm{I}})$ and $p_{1}(\mathbf{x}|\mathbf{x}_{1})=\mathcal{N}(\mathbf{x}_{1},\sigma^{2}{\bm{I}})$, with small $\sigma>0$. That is, the distribution $p_{0}(\mathbf{x}|\mathbf{x}_{1})$ corresponds to the noise distribution and the distribution $p_{1}(\mathbf{x}|\mathbf{x}_{1})$ is centered around the data point with small variance.
Then there exists a conditional vector field $u_{s}(\mathbf{x}|\mathbf{x}_{1})$ which generates $p_{s}(\mathbf{x}|\mathbf{x}_{1})$. Our goal is to learn the vector field with a neural network $\epsilon_{\theta}(\mathbf{x},s)$ which amounts to learning the generative process. This can be done by minimizing the flow matching objective:
$\begin{split}\mathcal{L}=&E_{s\sim\mathcal{U}(0,1),\mathbf{x}_{1}\sim\text{Data},\mathbf{x}\sim p_{s}(\mathbf{x}|\mathbf{x}_{0})}\Big{[}\\ &\lVert\epsilon_{\theta}(\mathbf{x},s)-u_{s}(\mathbf{x}_{s}|\mathbf{x}_{1})\rVert_{2}^{2}\Big{]}.\end{split}$ (7)
Going back to Equation 6 we notice that the two approaches have similarities. Flow matching differs in the path constructions and it learns the vector field directly, instead of learning the score, potentially offering a more stable alternative.
One choice for the noising function is transporting the values into noise as a linear function of transport time:
$\mathbf{x}_{s}=s\mathbf{x}_{1}+(1-(1-\sigma)s)\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}({\bm{0}},{\bm{I}}).$ (8)
The probability path is generated by the following conditional vector field which is available in closed-form:
$u_{s}(\mathbf{x}|\mathbf{x}_{1})=\frac{\mathbf{x}_{1}-(1-\sigma)\mathbf{x}}{1-(1-\sigma)s}.$ (9)
By learning the field $u_{s}(\mathbf{x}|\mathbf{x}_{1})$ with a neural network $\epsilon_{\theta}(\mathbf{x},s)$ we can sample new points by sampling an initial value $\mathbf{x}_{0}$ from the noise distribution $p_{0}$ and solve an ODE $0\mapsto 1$ to obtain the new sample $\mathbf{x}_{1}$.
2.4 Stochastic Interpolants (SI)
Stochastic interpolants [Albergo et al., 2023b] aim to model the dependent couplings between $(\mathbf{x}_{0},\mathbf{x}_{1})$ with their joint density $\rho(\mathbf{x}_{0},\mathbf{x}_{1})$, and establish a two-way generative SDEs mapping from one data distribution to another. The method constructs a straightforward stochastic mapping from $t=0$ to $t=1$ given the values at two ends $\mathbf{x}_{0}\sim\rho_{0}$ to $\mathbf{x}_{1}\sim\rho_{1}$, which provides a transport between two densities $\rho_{0}$ and $\rho_{1}$, while maintaining the dependency between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$.
$\mathbf{x}_{s}=\alpha(s)\mathbf{x}_{0}+\beta(s)\mathbf{x}_{1}+\gamma(s)\mathbf{z},\;s\in[0,1],\mathbf{z}\sim\mathcal{N}(\mathbf{0},I)$ (10)
where $\rho(s,\mathbf{x})$ is the marginal density of $\mathbf{x}_{s}$ at diffusion time $s$. Such a stochastic mapping is characterized by a pair of functions: velocity function $\mathbf{b}(s,\mathbf{x})$ and score function $\mathbf{s}(s,\mathbf{x})$:
$\mathbf{s}(s,\mathbf{x}):=\nabla\log\rho(s,\mathbf{x}),$ (11)
$\mathbf{b}(s,\mathbf{x}):=\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{z}}[\dot{\alpha}(s)\mathbf{x}_{0}+\dot{\beta}(s)\mathbf{x}_{1}+\dot{\gamma}(s)\mathbf{z}|\mathbf{x}_{s}=\mathbf{x}].$ (12)
$\mathbf{b}(s,\mathbf{x})$, $\rho(s,\mathbf{x})$, and $\mathbf{s}(s,\mathbf{x})$ satisfy the equality below,
$\partial_{t}\rho(s,\mathbf{x})+\nabla\cdot(\mathbf{b}(s,\mathbf{x})\rho(s,\mathbf{x}))=0$ (13)
$\mathbf{s}(s,\mathbf{x})=-\gamma^{-1}(s)\mathbb{E}_{\mathbf{z}}[\mathbf{z}|\mathbf{x}_{t}=\mathbf{x}],$ (14)
where $\alpha(s)$ and $\beta(s)$ schedule the deterministic interpolant. We set $\alpha(0)=1,\alpha(1)=0,\beta(0)=0,\beta(1)=1$. $\gamma(s)$ schedules the variance of the stochastic component $\mathbf{z}$. We set $\gamma(0)=\gamma(1)=0$, so the two ends of the interpolant are fixed at $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. Figure 1 shows one example of the interpolant schedule, where $\alpha(s)=\sqrt{1-\gamma^{2}(s)}\cos(\frac{1}{2}\pi s)$, $\beta(s)=\sqrt{1-\gamma^{2}(s)}\sin(\frac{1}{2}\pi s)$, $\gamma(s)=\sqrt{2s(1-s)}$.
Figure 1: $\alpha(s),\beta(s),\gamma(s)$, the schedules of stochastic interpolants.
The velocity function $\mathbf{b}(s,\mathbf{x})$ and the score function $\mathbf{s}(s,\mathbf{x})$ can be modeled by a rich family of functions, such as deep neural networks. The model is trained to match the above equality by minimizing the mean squared error loss functions,
$\mathcal{L}_{b}=$ $\int_{0}^{1}\mathbb{E}\Big{[}\frac{1}{2}\|\hat{\mathbf{b}}(s,\mathbf{x}_{s})\|^{2}$ (15)
$-\big{(}\dot{\alpha}(s)\mathbf{x}_{0}+\dot{\beta}(s)\mathbf{x}_{1}+\dot{\gamma}(s)\mathbf{z}\big{)}^{T}\hat{\mathbf{b}}(s,\mathbf{x}_{s})\Big{]}{\mathrm{d}}s$
$\mathcal{L}_{s}=$ $\int_{0}^{1}\mathbb{E}\Big{[}\frac{1}{2}\|\hat{\mathbf{s}}(s,\mathbf{x}_{s})\|^{2}+\gamma^{-1}\mathbf{z}^{T}\hat{\mathbf{s}}(s,\mathbf{x}_{s})\Big{]}{\mathrm{d}}s.$ (16)
More details of training will be shown in section 4.
During inference, usually, one side of the diffusion trajectory at $t=0$ or $t=1$ is given, the goal is to infer the sample distribution on the other side. The interpolant in equation 10 results in elegant forward and backward SDEs and corresponding Fokker-Planck equations, which offer convenient tools for inference. The SDEs are composed of $\mathbf{b}(s,\mathbf{x}_{s})$ and $\mathbf{s}(s,\mathbf{x}_{s})$, which are learned from the data. For any $\epsilon(s)\geq 0$, define the forward and backward SDEs
${\mathrm{d}}\mathbf{x}_{s}=$ $[\mathbf{b}(s,\mathbf{x})+\epsilon(s)\mathbf{s}(s,\mathbf{x})]{\mathrm{d}}s+\sqrt{2\epsilon(s)}{\mathrm{d}}\mathbf{w}_{s}$ (17)
${\mathrm{d}}\mathbf{x}_{s}=$ $[\mathbf{b}(s,\mathbf{x})-\epsilon(s)\mathbf{s}(s,\mathbf{x})]{\mathrm{d}}s+\sqrt{2\epsilon(s)}{\mathrm{d}}\mathbf{w}^{\mathrm{B}}_{s},$ (18)
where $\mathbf{w}^{\mathrm{B}}_{s}$ is the backward Brownian motion. The SDEs satisfy the forward and backward Fokker-Plank equations,
$\partial_{s}\rho+\nabla\cdot(\mathbf{b}_{\mathrm{F}}\rho)=\epsilon(s)\Delta\rho,\rho(0)=\rho_{0}$ (19)
$\partial_{s}\rho+\nabla\cdot(\mathbf{b}_{\mathrm{B}}\rho)=-\epsilon(s)\Delta\rho,\rho(1)=\rho_{1}.$ (20)
These properties imply that one can draw samples from the conditional density $\rho(\mathbf{x}_{1}|\mathbf{x}_{0})$ following the forward SDE in equation 17 starting from $\mathbf{x_{0}}$ at $s=0$. It can also draw samples from the joint density $\rho(\mathbf{x}_{0},\mathbf{x}_{1})$ by initially drawing a sample $\mathbf{x}_{0}\sim\rho_{0}$ (if feasible, for example, pick one sample from the dataset), then using the forward SDE to generate a samples $\mathbf{x}_{1}$ at $s=1$. The method guarantees that $\mathbf{x}_{1}$ follows marginal distribution $\rho_{1}$ and the sample pair $(\mathbf{x}_{0},\mathbf{x}_{1})$ satisfies the joint density $\rho(\mathbf{x}_{0},\mathbf{x}_{1})$. Drawing samples using the backward SDE is similar: one can draw samples from $\rho(\mathbf{x}_{0}|\mathbf{x}_{1})$ and the joint density $\rho(\mathbf{x}_{0},\mathbf{x}_{1})$ as well. Details of inference will be shown in section 4.
Figure 2: Stochastic interpolants for time series prediction using forward SDE in equation 22. Yu: Change rrn input index.
3 Conditional generation with extra features
All the aforementioned methods can be adapted for conditional generation with additional features. The conditions may range from simple categorical values [Song et al., 2021] to complex prompts involving multiple data types, including partial observations of a sample’s entries (e.g., image inpainting, time series imputation) [Tashiro et al., 2021, Song et al., 2021], images [Zheng et al., 2023, Rombach et al., 2022], text [Rombach et al., 2022, Zhang et al., ], etc. A commonly employed technique to handle diverse conditions is to integrate condition information through feature embedding, where the embedding is injected into various layers of neural networks [Song et al., 2021, Rombach et al., 2022]. For instance, conditional SGM can be trained with
$\mathcal{L}_{\mathrm{cond}}=$ $\mathbb{E}_{s\sim\mathcal{U}(0,1),(\mathbf{x}_{0},\xi)\sim\text{Data},\mathbf{x}_{s}\sim p(\mathbf{x}_{s}|\mathbf{x}_{0})}\Big{[}$ (21)
$\lVert\epsilon_{\theta}(s,\mathbf{x}_{s},\xi)-\nabla_{\mathbf{x}_{s}}\log p(\mathbf{x}_{s}|\mathbf{x}_{0})\rVert_{2}^{2}\Big{]}.$
where the data is given by pairs of a sample $\mathbf{x}_{0}$ and the corresponding condition $\xi$. This simple scheme approach showcases its effectiveness in various tasks, achieving state-of-the-art performance [Rombach et al., 2022, Zhang et al., ].
Figure 3: Examples of model generated samples for synthetic two-dimensional datasets.
Likewise, SI can be expanded for conditional generation by substituting the velocity function and score function with $\mathbf{b}(\mathbf{x}_{s},s,\xi)$ and $\mathbf{s}(\mathbf{x}_{s},s,\xi)$ [Albergo et al., 2023b]. The model is trained using samples of tuples $(\mathbf{x}_{0},\mathbf{x}_{1},\xi)$, where $\xi$ is the extra condition feature. Consequently, the inference using forward or backward SDEs becomes
${\mathrm{d}}\mathbf{x}_{s}=$ $[\mathbf{b}(s,\mathbf{x}_{s},\xi)+\epsilon(s)\mathbf{s}(s,\mathbf{x}_{s},\xi)]{\mathrm{d}}s+\sqrt{2\epsilon(s)}{\mathrm{d}}\mathbf{w}_{s}$ (22)
${\mathrm{d}}\mathbf{x}_{s}=$ $[\mathbf{b}(s,\mathbf{x}_{s},\xi)-\epsilon(s)\mathbf{s}(s,\mathbf{x}_{s},\xi)]{\mathrm{d}}s+\sqrt{2\epsilon(s)}{\mathrm{d}}\mathbf{w}^{\mathrm{B}}_{s},$ (23)
where both velocity and score functions depend on the condition $\xi$. The loss functions are similar to equation 15 and equation 16.
Regarding the time series prediction task, we will encode a large context window as the conditional information, and the prediction or generation of future time points will rely on such a conditional generation mechanism.
Next, we demonstrate that the probability distribution of ${\mathrm{x}}_{t}$ as simulated by equation 24, results in a dynamic density function. This density serves as a solution to a transport equation 25, which smoothly transitions between $\rho_{0}$ and $\rho_{1}$.
Algorithm 1 Training algorithm. $i$ is the sample index.
  Input: Sample 3-tuples $(\mathbf{x}_{t_{i}+1},\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i}-P:t_{i}-1})$. Interpolant schedules: $\alpha(s),\beta(s),\gamma(s)$. Models: velocity $\hat{\mathbf{b}}$, score $\hat{\mathbf{s}}$, $\mathrm{RNN}$.
  for iteration $=1$ to total iterations do
     $s_{i}\sim\mathrm{Beta}(0.1,0.1)$. $\mathbf{x}_{s,i}=\alpha(s_{i})\mathbf{x}_{t_{i}}+\beta(s_{i})\mathbf{x}_{t_{i}+1}+\gamma(s_{i})\mathbf{z}_{i}$ $\mathbf{h}_{i}=\mathrm{RNN}(\mathbf{x}_{t_{i}-P:t_{i}-1})$
$\mathcal{L}_{b}=\sum_{i=1}^{\text{batch size}}\frac{1}{p_{\mathrm{Beta}}(s_{i})}\Big{[}\frac{1}{2}\|\hat{\mathbf{b}}(s_{i},\mathbf{x}_{s,i},\mathbf{h}_{i})\|^{2}$
$-\big{(}\dot{\alpha}(s_{i})\mathbf{x}_{t_{i}}+\dot{\beta}(s_{i})\mathbf{x}_{t_{i}+1}+\dot{\gamma}(s_{i})\mathbf{z_{i}}\big{)}^{T}\hat{\mathbf{b}}(s_{i},\mathbf{x}_{s,i},\mathbf{h}_{i})\Big{]}$
$\mathcal{L}_{s}=\sum_{i=1}^{\text{batch size}}\frac{1}{p_{\mathrm{Beta}}(s_{i})}\Big{[}\frac{1}{2}\|\hat{\mathbf{s}_{i}}(s_{i},\mathbf{x}_{s,i})\|^{2}$
$+\gamma^{-1}\mathbf{z_{i}}^{T}\hat{\mathbf{s}_{i}}(s_{i},\mathbf{x}_{s,i})\Big{]}$
Perform back-propagation by minimizing $\mathcal{L}_{b}$ and $\mathcal{L}_{s}$.
  end for
Theorem 1.
(Extension of Stochastic Interpolants to Arbitrary Joint Distributions). Let $\rho_{01}$ be the joint distribution $(x_{0},x_{1})\sim\rho_{01}$ and let the stochastic interpolant be
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{1}+\gamma_{t}\mathbf{z},$ (24)
where $\alpha_{0}=\beta_{1}=1$, $\alpha_{1}=\beta_{0}=\gamma_{0}=\gamma_{1}=0$, and $\alpha_{t}^{2}+\beta_{t}^{2}+\gamma_{t}^{2}>0$ for all $t\in[0,1]$. We define $\rho_{t}$ to be the time-dependent density of $\mathbf{x}_{t}$, which satisfies the boundary conditions at $t=0,1$ and the transport equation follows that
$\dot{\rho}_{t}+\nabla\cdot(\mathbf{b}_{t}\rho_{t})=0$ (25)
for all $t\in[0,1]$ with the velocity defined as
$\mathbf{b}_{t}(\mathbf{x}|\xi)=\mathbb{E}\left[\dot{\alpha_{t}}\mathbf{x}_{0}+\dot{\beta_{t}}\mathbf{x}_{1}+\dot{\gamma}_{t}\,\mathbf{z}|\mathbf{x}_{t}=\mathbf{x},\xi\right],$ (26)
where the expectation is based on the density $\rho_{01}$ given $\mathbf{x}_{s}=\mathbf{x}$ and the extra information $\xi$.
The score function follows the relation such that
$\nabla\log\rho_{t}(\mathbf{x})=-\gamma^{-1}_{t}\mathbb{E}\left[\mathbf{z}|\mathbf{x}_{t}=\mathbf{x},\xi\right].$
The proof is in a spirit similar to Theorem 2 in [Albergo et al., 2023b] and detailed in section B. The key difference is that we consider a continuous-time interpretation and avoid using characteristic functions, which makes the analysis more friendly to users. Additionally, the score function $\nabla\log\rho_{t}(x)$ is optimized in a simple quadratic objective function as indicated in Theorem 2 in the Appendix.
4 Stochastic interpolants for time series prediction
We formulate time series prediction tasks through the conditional probability $p(\mathbf{x}_{t+1}|\mathbf{x}_{t-P:t})$. The model diagram is illustrated in Figure 2. Here, $\mathbf{x}_{t}\in\mathbb{R}^{D}$ represents the multivariate time series at time $t$ with $D$ dimensions, $\mathbf{x}_{t+1}$ is the prediction target, and $\mathbf{x}_{t-P:t}$ is the context window, where $P$ denotes the length of the context window.
For this problem, we employ the conditional Stochastic Interpolants (SI) method as follows. In the training phase, the generative model learns the joint distribution $p(\mathbf{x}_{t+1},\mathbf{x}_{t}|\mathbf{x}_{t-P:t-1})$ of the pair $(\mathbf{x}_{t+1},\mathbf{x}_{t})$ given the past observations $\mathbf{x}_{t-P:t-1}$. where $\mathbf{x}_{t}\sim\rho_{0}$ and $\mathbf{x}_{t+1}\sim\rho_{1}$ for all $t$, so the marginal distributions are equal $\rho_{0}=\rho_{1}$. The model aims to learn the coupling relation between $\mathbf{x}_{t+1}$ and $\mathbf{x}_{t}$ conditioning on the context $\mathbf{x}_{t-P:t-1}$. This is achieved by training the conditional velocity and score functions in equation 22.
Algorithm 2 Inference algorithm
  Input: Sample 2-tuples $(\mathbf{x}_{t},\mathbf{x}_{t-P:t-1})$. Trained models: Velocity $\hat{\mathbf{b}}$, score $\hat{\mathbf{s}}$, $\mathrm{RNN}$. Diffusion variance $\epsilon$. Set $\tilde{\mathbf{x}}_{0}=\mathbf{x}_{t}$. $\mathbf{h}=\mathrm{RNN}(\mathbf{x}_{t-P:t-1})$. Run SDE integral for $s\in[0,1]$ following
${\mathrm{d}}\tilde{\mathbf{x}}_{s}=[\mathbf{b}(s,\tilde{\mathbf{x}}_{s},\mathbf{h})+\epsilon\mathbf{s}(s,\tilde{\mathbf{x}}_{s},\mathbf{h})]{\mathrm{d}}s+\sqrt{2\epsilon}{\mathrm{d}}\mathbf{w}_{s}$
  Output: $\tilde{\mathbf{x}}_{1}$ as prediction of $\mathbf{x}_{t+1}$.
8gaussians Circles Moons Rings Swissroll
DDPM (cosine) 2.58 0.20 0.20 0.12 0.24
DDPM (linear) 0.70 0.18 0.12 0.11 0.14
SGM 1.10 0.30 0.35 0.32 0.14
FM 0.58 0.10 0.11 0.09 0.15
SI (quad, linear) 0.52 0.15 0.32 0.12 0.16
SI (sqrt, linear) 0.59 0.29 0.51 0.22 0.37
SI (sqrt, trig) 0.75 0.25 0.50 0.48 0.36
SI (trig, linear) 0.52 0.13 0.29 0.21 0.16
Table 1: Wasserstein distance between the generated samples and true data.
As the sample spaces of $\rho_{0}$ and $\rho_{1}$ must be the same, the generative model can not directly map the whole context window $\mathbf{x}_{t-P:t}$ to the target $\mathbf{x}_{t+1}$ due to different vector sizes. Instead, a recurrent neural network is used to encode the context $\mathbf{x}_{t-P:t-1}$ into a history prompt $\mathbf{h}_{t}$. Subsequently, the score function and velocity function perform conditional generation diffusing from $\mathbf{x}_{t}$ with the condition input $\mathbf{h}_{t}$ following equation 22.
The training dataset consists of tuple $(\mathbf{x}_{t+1},\mathbf{x}_{t},\mathbf{x}_{t-P:t-1})$. It is worth noting that the loss values become larger when $s$ is close to two ends. To address this, importance sampling is leveraged to better handle the integral over diffusion time in the loss functions equation 15 and equation 16 to stabilize the training, where we use Beta distribution for our proposal distribution. The algorithm is outlined in Algorithm 1. Additional details can be found in Appendix C.
In the inference phase, the RNN first encodes the context $\mathbf{x}_{t-P:t-1}$ into the history prompt $\mathbf{h}_{t}$, then SI transports the context vector $\mathbf{x}_{t}$ to the target distribution with the condition $\mathbf{h}_{t}$, following the forward SDE. Regarding the multiple steps prediction, we recursively run the step-by-step prediction.
Figure 4: Forecast paths for SI on Solar dataset showing median prediction, 50th and 90th confidence intervals calculated from model samples, on 6/ 13761376\;/\;1376 / 137 variate dimensions.
5 Experiments
We first verify the method on synthetic datasets, and then apply it to the time series forecasting tasks with real data.
Figure 5: Time-Grad model for conditional time series prediction.
Baseline models such as DDPM, SGM, FM, and SI all involve modeling field functions, where the inputs are the state vector (in the same space of the data samples), diffusion time, and condition embedding, and the output is the generated sample. The field functions correspond to the noise prediction function in DDPM; the score function in SGM; the vector field in FM; and the velocity and score functions in SI. To make a fair comparison between these models, we use the same neural networks for these models. Details of the models are discussed in Appendix C. Our code will be publicly available111github_place_holder We will publish the code for the camera-ready version..
Exchange rate Solar Traffic Wiki
Vec-LSTM 0.008$\pm$0.001 0.391$\pm$0.017 0.087$\pm$0.041 0.133$\pm$0.002
DDPM 0.009$\pm$0.004 0.359$\pm$0.061 0.058$\pm$0.014 0.084$\pm$0.023
FM 0.009$\pm$0.001 0.419$\pm$0.027 0.038$\pm$0.002 64.256$\pm$62.596
SGM 0.008$\pm$0.002 0.364$\pm$0.029 0.071$\pm$0.05 0.108$\pm$0.026
SI 0.007$\pm$0.001 0.359$\pm$0.06 0.083$\pm$0.005 0.080$\pm$0.007
Table 2: CRPS-sum metric on multivariate probabilistic forecasting. A smaller number indicates better performance.
5.1 Synthetic datasets
We synthesize several two-dimensional datasets with regular patterns, such as Circles, Moons, etc. Details can be found in [Chen et al., 2021, Lipman et al., 2022] and their published code repositories. Model introduced in section 2 are compared to the SI as baselines. For diffusion-like models, we implement DPPM with a linear or cosine noise scheduler. We explore on synthetic datasets to determine a good range of hyperparameters which will be used in later time series experiments. This experiment is used to investigate the properties with respect to the varying data sizes, model sizes, and training lengths.
To fairly compare the generation quality, all models are assigned to generate data in the same setting by mapping from standard Gaussian to the target distribution. The neural networks and hyperparameters are also set as the same, such as batch size, training epochs, etc. The generated samples from different methods are shown in Figure 3. Table 1 measures the sample quality with Wasserstein distance [Ramdas et al., 2017]. It shows that all the models can capture the true distribution. The same holds when we use different metrics such as Sliced Wasserstein Distance (SWD) [Rabin et al., 2012] and Maximum Mean Discrepancy (MMD) [Gretton et al., 2012].
We also test-out different schedulers for a stochastic interpolant model. For example, “SI (sqrt, linear)” means we use square root gamma-function $\gamma(s)=\sqrt{2s(1-s))}$ and a linear interpolant. Other gamma-functions that we consider are quad: $\gamma(s)=t(1-s)$, and trig: $\gamma(s)=\sin(\pi t)^{2}$. We show that most of the gamma-interpolant function combinations achieve good results modeling the target distribution.
5.2 Multivariate probabilistic forecasting
In this section, we will empirically verify that: 1) SI is a suitable generative module for the prediction compared with other baselines with different generative methods under the same framework; 2) the whole framework can achieve competitive performance in time series forecasting.
Models.
The baseline models include DDPM, SGM, and FM-based generative models adopted for step-by-step (autoregressive) prediction. DDPM and SGM-based models can only generate samples by transporting Gaussian noise distribution to data distribution. So we modify the framework by replacing the context time point $\mathbf{x}_{t}$ with Gaussian noise, as shown in Figure 5. Flow matching can easily fit into this framework by replacing the denoising objective with the flow matching objective. The modified framework is shown in Figure 2. We model the map from the previous time series observation to the next (forecasted) value. We argue this is a more natural choice than mapping from noise for each time series prediction step. Recent Vec-LSTM from [Salinas et al., 2019] is compared as a pure recurrent neural network model whose probabilistic layer is a multivariate Gaussian.
Setup.
The real-world time series datasets include Solar [Lai et al., 2018], Exchange [Lai et al., 2018], Traffic [^2] https://archive.ics.uci.edu/ml/datasets/PEMS-SF, and Wiki [^3] https://github.com/mbohlkeschneider/gluon-ts/tree/mv_release/datasets which have been commonly used for the probabilistic forecasting tasks. We follow the preprocessing steps as in [Salinas et al., 2019]. The probabilistic forecasting is evaluated by Continuous Ranked Probability Score (CRPS-sum) [Koochali et al., 2022], normalized root mean square error via the median of the samples (NRMSE), and point-metrics normalized deviance (ND). The metrics calculation is provided by gluonts package [Alexandrov et al., 2019]. In all of the cases smaller values indicate better performance.
Results.
The results for CRPS-sum are shown in Table 2. The results for other metrics are consistent with CRPS-sum and are shown in Tables 4 and 5, in Appendix C. We outperform or match other models on three out of four datasets, only on Traffic FM model achieves better performance. Note that on Wiki data FM cannot capture the data distribution. We ran a search over flow matching hyperparameters without being able to get satisfying results. Therefore, we conclude that stochastic interpolants are a strong candidate for conditional generation, in particular for multivariate probabilistic forecasting. By comparing to the RNN-based model Vec-LSTM, our model and other baselines such as SGM and DDPM get better performance, which implies that carefully model the probability distribution is critical for large dimension time series prediction. Figure 4 demonstrates the quality of the forecast on Solar dataset. We can see that our model can make precise prediction and capture the uncertainty, even when the scale of the different dimensions varies a lot.
6 Conclusions
This study presents an innovative method that effectively merges the computational efficiency of recurrent neural networks with the high-quality probabilistic modeling of the diffusion model, specifically applied to probabilistic time series forecasting. Grounded in stochastic interpolants and an expanded conditional generation framework featuring control features, the method undergoes empirical evaluation on both synthetic and real datasets, showcasing its compelling performance.
Appendix A Related Works
A plethora of papers focus on auto-regression models, particularly transformer-based models. For a more comprehensive review, we refer to Wen et al. [2023], Chen et al. [2023c]. While our work does not aim to replace RNN- or transformer-based architectures, we emphasize that one of the main motivations behind our work is to develop a probabilistic module building upon these recent advancements. Due to limited resources, we did not extensively explore all underlying temporal architectures but instead selected relatively simpler models as defaults.
The authors were aware of other diffusion-based probabilistic models, as highlighted in the introduction. Unlike our lightweight model, which models the transition between adjacent time points, these selected works model the entire time window, requiring both high memory and computational complexity. With our computation budget restricted to a 32 GB GPU device, effectively training these diffusion models on large datasets with hundreds of features is challenging.
Additionally, several relevant works are related to our idea. For instance, Rasul et al. [2021b] incorporates the DDPM structure, aligning with our DDPM baseline structure. During inference, the prediction diffuses from pure noise to the target distribution. TimeDiff Shishkov [2023] introduces two modifications to the established diffusion model: during training it mixes target and context data and it adds an AR model for more precise initial prediction. Both of these can be incorporated into our model as well.
The existing probabilistic forecasters model the distribution of the next value from scratch, meaning they start with normal distribution in the case of normalizing flows and diffusion models or output parametric distribution in the case of transformers and deep AR. We propose modeling the transformation between the previously observed value and the next value we want to predict. We believe this is a more natural way to forecast which can be seen from requiring fewer solver steps to reach the target distribution.
The second row (DDPM) in Table 2 is an exact implementation of Rasul et al. [2021b]. The results might be different due to slightly different training setups but all the models share the same training parameters so the rank should remain the same. We also include ND-sum and NRMSE-sum in the appendix for completeness.
Discussion of Vec-LSTM baseline
In terms of the neural network architecture, we use similar architecture for the LSTM encoder. But to be clear, Vec-LSTM Salinas et al. [2019] and our SI framework are not exactly the same mainly due to different ways of probabilistic modeling. Vec-LSTM considers the multivariate Gaussian distribution for the time points, where the mean and covariance matrices are modeled using separate LSTMs. Especially, the covariance matrix is modeled through a low-dimensional structure $\bm{\Sigma}(\bm{h}_{t})=\bm{D}_{t}(\bm{h}_{t})+\bm{V}_{t}(\bm{h}_{t})\bm{V}_{t}(\bm{h}_{t})^{\intercal}$, where $\bm{h}_{t}$ is the latent variable from LSTM. The SI framework does not explicitly model the output distribution in any parametric format. Instead, the latent variable from RNN output is used as the condition variable to guide the diffusion model in Eq.22. Thus, the architectures of RNNs in the two frameworks are not quite strictly comparable.
Appendix B Proof: Conditional Stochastic Interpolant
The proof is in spirit similar to Theorem 2 in Albergo et al. [2023b]. The key difference is that we consider a continuous-time interpretation, which makes the analysis more friendly to users.
Proof [Proof of Theorem 1]
Given the conditional information $\xi$ and ${\mathrm{x}}_{s}={\mathrm{x}}$ simulated from equation 24, the conditional stochastic interpolant for equation 24 follows that
$\mathbb{E}[{\mathrm{x}}_{t}|{\mathrm{x}}_{s}={\mathrm{x}},\xi]=\mathbb{E}[\alpha_{t}{\mathrm{x}}_{0}+\beta_{t}{\mathrm{x}}_{1}+\gamma_{t}z|{\mathrm{x}}_{s}={\mathrm{x}},\xi],$ (27)
where the expectation takes over the density for $({\mathrm{x}}_{0},{\mathrm{x}}_{1})\sim\rho({\mathrm{x}}_{0},{\mathrm{x}}_{1}|\xi)$, $\xi\sim\eta(\xi)$, and $z\sim\mathrm{N}(0,\mathrm{I})$.
We next show equation 27 is a solution of a stochastic differential equation as follows
${\textnormal{d}}\mathbb{E}[{\mathrm{x}}_{t}|{\mathrm{x}}_{s}={\mathrm{x}},\xi]$ $={\mathbf{f}}_{t}({\mathrm{x}}){\mathrm{d}}t+\sigma_{t}{\textnormal{d}}{\mathrm{w}}_{t},$ (28)
where ${\mathbf{f}}_{t}({\mathrm{x}})=\mathbb{E}[\dot{\alpha_{t}}{\mathrm{x}}_{0}+\dot{\beta_{t}}{\mathrm{x}}_{1}|{\mathrm{x}}_{s}={\mathrm{x}},\xi]$ and $\sigma_{t}=\sqrt{2{\gamma_{t}}\dot{\gamma_{t}}}$.
To prove the above argument, we proceed to verify the drift and diffusion terms respectively:
  • Drift: It is straightforward to verify the drift ${\mathbf{f}}_{t}$ by taking the gradient of the conditional expectation $\mathbb{E}[\alpha_{t}{\mathrm{x}}_{0}+\beta_{t}{\mathrm{x}}_{1}|{\mathrm{x}}_{s}={\mathrm{x}},\xi]$ with respect to $t$.
  • Diffusion: For the diffusion term, the proof hinges on showing $\sigma_{t}=\sqrt{2{\gamma_{t}}\dot{\gamma_{t}}}$, which boils down to prove the stochastic calculus follows that $\int_{0}^{t}\sqrt{2{\gamma_{s}}\dot{\gamma_{s}}}{\textnormal{d}}{\mathrm{w}}_{s}=\gamma_{t}z$. Note that $\mathbb{E}[\int_{0}^{t}\sqrt{2{\gamma_{s}}\dot{\gamma_{s}}}{\textnormal{d}}{\mathrm{w}}_{s}]=0$. Invoking the Itô isometry, we have $\text{Var}(\int_{0}^{t}\sqrt{2{\gamma_{s}}\dot{\gamma_{s}}}{\textnormal{d}}{\mathrm{w}}_{s})=\int_{0}^{t}2{\gamma_{s}}\dot{\gamma_{s}}{\mathrm{d}}s=\int_{0}^{t}({\gamma_{s}}^{2})^{\prime}{\mathrm{d}}s=\gamma_{t}^{2}$ (given $\gamma_{0}=0$). In other words, $\int_{0}^{t}\sqrt{2{\gamma_{s}}\dot{\gamma_{s}}}{\textnormal{d}}{\mathrm{w}}_{s}$ is a normal random variable with mean 0 and variable $\gamma_{t}^{2}$, which proves that equation 27 is a solution of the stochastic differential equation 28.
Define $\Sigma_{t}=2{\gamma_{t}}\dot{\gamma_{t}}$, we know the Fokker-Planck equation associated with equation 28 follows that
$\begin{split}0&=\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot\bigg{(}{\mathbf{f}}_{t}\rho_{t}-\frac{1}{2}\Sigma_{t}\nabla\rho_{t}\bigg{)}\\ &=\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot\bigg{(}\bigg{(}{\mathbf{f}}_{t}-\frac{1}{2}\Sigma_{t}\nabla\log\rho_{t}\bigg{)}\rho_{t}\bigg{)}\\ &=\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot\bigg{(}\bigg{(}\mathbb{E}[\dot{\alpha_{t}}{\mathrm{x}}_{0}+\dot{\beta_{t}}{\mathrm{x}}_{1}|{\mathrm{x}}_{s}={\mathrm{x}},\xi]-{\gamma_{t}}\dot{\gamma_{t}}\nabla\log\rho_{t}\bigg{)}\rho_{t}\bigg{)}\\ &=\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot\big{(}b_{t|s}({\mathrm{x}},\xi)\rho_{t}\big{)},\end{split}$ (29)
where $b_{t|s}({\mathrm{x}}|\xi)=\mathbb{E}[\dot{\alpha_{t}}{\mathrm{x}}_{0}+\dot{\beta_{t}}{\mathrm{x}}_{1}-{\gamma_{t}}\dot{\gamma_{t}}\nabla\log\rho_{t}|{\mathrm{x}}_{s}={\mathrm{x}},\xi]$.
Further setting $s=t$ and rewrite $b_{t}\equiv b_{t|t}$, we have $b_{t}({\mathrm{x}}|\xi)=\mathbb{E}[\dot{\alpha_{t}}{\mathrm{x}}_{0}+\dot{\beta_{t}}{\mathrm{x}}_{1}-{\gamma_{t}}\dot{\gamma_{t}}\nabla\log\rho_{t}|{\mathrm{x}}_{t}={\mathrm{x}},\xi]$
Further define $g^{(i)}_{t}({\mathrm{x}}|\xi)=\mathbb{E}[{\mathrm{x}}_{i}|{\mathrm{x}}_{t}={\mathrm{x}},\xi]$, where $i\in\{0,1\}$ and $g^{(z)}_{t}({\mathrm{x}}|\xi)=\mathbb{E}[z|{\mathrm{x}}_{t}={\mathrm{x}},\xi]$. We have that
$b_{t}({\mathrm{x}}|\xi)$ $=\mathbb{E}[\dot{\alpha_{t}}{\mathrm{x}}_{0}+\dot{\beta_{t}}{\mathrm{x}}_{1}-{\gamma_{t}}\dot{\gamma_{t}}\nabla\log\rho_{t}|{\mathrm{x}}_{t}={\mathrm{x}},\xi]$
$=\dot{\alpha_{t}}g^{(0)}+\dot{\beta_{t}}g^{(1)}+\dot{\gamma_{t}}g^{(z)}$
$=\mathbb{E}[\dot{\alpha_{t}}{\mathrm{x}}_{0}+\dot{\beta_{t}}{\mathrm{x}}_{1}+\dot{\gamma_{t}}z|{\mathrm{x}}_{t}={\mathrm{x}},\xi],$
where the first equality follows by equation 29 and the last one follows by taking derivative to equation 27 w.r.t. the time $t$.
We also observe that $\nabla\log\rho_{t}=-\gamma_{t}^{-1}\mathbb{E}[z|{\mathrm{x}}_{t}={\mathrm{x}}]$.
Theorem 2.
The loss functions used for estimating the vector field follow that
$L_{i}(\hat{g}^{(i)})=\int_{0}^{1}\mathbb{E}[|\hat{g}^{(i)}|^{2}-2{\mathrm{x}}_{i}\cdot\hat{g}^{(i)}]{\mathrm{d}}t,$
where $i\in\{0,1,z\}$, the expectation takes over the density for $({\mathrm{x}}_{0},{\mathrm{x}}_{1})\sim\rho({\mathrm{x}}_{0},{\mathrm{x}}_{1}|\xi)$, $\xi\sim\eta(\xi)$, and $z\sim\mathrm{N}(0,\mathrm{I})$.
Proof  To show the loss is effective to estimate $g^{(0)}$, $g^{(1)}$, and $g^{(z)}$. It suffices to show
$L_{0}(\hat{g}^{(0)})$ $=\int_{0}^{1}\mathbb{E}[|\hat{g}^{(0)}|^{2}-2{\mathrm{x}}_{0}\cdot\hat{g}^{(0)}]{\mathrm{d}}t,$
$=\int_{0}^{1}\int_{\mathrm{R}^{d}}\bigg{[}|\hat{g}^{(0)}|^{2}-2\mathbb{E}[{\mathrm{x}}_{0}|{\mathrm{x}}_{t}={\mathrm{x}},\xi]\cdot\hat{g}^{(0)}\bigg{]}{\mathrm{d}}{\mathrm{x}}{\mathrm{d}}t,$
$=\int_{0}^{1}\int_{\mathrm{R}^{d}}\bigg{[}|\hat{g}^{(0)}|^{2}-2g^{(0)}\cdot\hat{g}^{(0)}\bigg{]}{\mathrm{d}}{\mathrm{x}}{\mathrm{d}}t,$
where the last equality follows by definition. The unique minimizer is attainable by setting $\hat{g}^{(0)}=g^{(0)}$.
The proof of $g^{(1)}$ and $g^{(z)}$ follows a similar fashion.
Appendix C Experiment details
C.1 Time series data
The time series datasets include: SolarLai et al. [2018], ExchangeLai et al. [2018], Traffic [^4] https://archive.ics.uci.edu/ml/datasets/PEMS-SF, and Wikipedia5555https://github.com/mbohlkeschneider/gluon-ts/tree/mv_release/datasets. We follow the preprocessing steps as in Salinas et al. [2019]. Details of the datasets are listed in Table 3.
Table 3: Properties of the datasets.
Datasets Dimension Frequency Total time points Context length
Exchange 8 Daily 6,071 30
Solar 137 Hourly 7,009 24
Traffic 963 Hourly 4,001 24
Wiki 2000 Daily 792 30
The probabilistic forecasting is evaluated by Continuous Ranked Probability Score (CRPS-sum) Koochali et al. [2022], normalized root mean square error via the median of the samples (NRMSE), and point-metrics normalized deviance (ND). The metrics calculation is provided by gluonts package Alexandrov et al. [2019] by calling module gluonts.evaluation.MultivariateEvaluator.
C.2 Models and hyperperameters
Baseline models such as DDPM, SGM, FM, and SI all involve modeling field functions, where the inputs are the state vector (in the same space of the data samples), diffusion time, and condition embedding, and the output is the generated sample. The field functions correspond to the "noise prediction" function in DDPM; the score function in SGM; the vector field in FM; the velocity and score functions in SI. To make fair comparison between these models, we use the same neural networks for these models.
In the synthetic datasets experiments, we model the field functions with a 4-layer ResNet, each layer has 256 intermediate dimensions. The batch size is 10,000 for all models and a model is trained with 20,000 iterations. The learning rate is $10^{-3}$.
In the time series forecasting experiments, the RNN for the history encoder has 1 layer and 128 latent dimension; The field function is modeled with Unet-like structure Ronneberger et al. [2015] with 8 residual blocks, and each block has 64 dimensions. To stabilize the training, we also use paired-sampling for the stochastic interpolants introduced by [Albergo et al., 2023a, Appendix C].
$\mathbf{x}_{s}=$ $\alpha(s)\mathbf{x}_{0}+\beta(s)\mathbf{x}_{1}+\gamma(s)\mathbf{z}$
$\mathbf{x}_{s}^{\prime}=$ $\alpha(s)\mathbf{x}_{0}+\beta(s)\mathbf{x}_{1}+\gamma(s)(-\mathbf{z})$
$s\in[0,1],\mathbf{z}\sim\mathcal{N}(\mathbf{0},I)$
The baseline models are trained with 200 epochs and 64 batch size with learning rate $10^{-3}$. The SI model is trained with 100 epochs and 128 batch size with learning rate $10^{-4}$. We find if the learning rate is too large, SI may not converge properly. All these model implementation details can be found in our code repository at www:github_place_holder (we will make it publicly available after acceptance).
C.3 Importance sampling
The loss functions for training the velocity and score functions are
[…]
Both loss functions involve the integral over diffusion time $s\in[0,1]$ in the form of
[…]
However, the loss values $l(s)$ has very large variance, especially when $s$ is near 00 or $1$. Figure 6 shows an example of the distribution of $l(s)$ across multiple $s$. The large variance slows down the convergence of training. To overcome this issue, we apply importance sampling, similar technique used by [Song et al., 2021, Sec. 5.1 ], to stabilize the training. In stead of drawing diffusion time from uniform distribution, importance sampling considers,
[…]
Ideally, one wants to keep $\frac{1}{\tilde{q}(s_{i})}l(s_{i})$ as constant as possible such that the variance of the estimation is minimum. The loss value $l(s)$ is very large when $s$ is close to 00 or $1$, and $l(s)$ is relatively flat in the middle, and the domain of $s$ is $[0,1]$, so we choose Beta distribution $\mathrm{Beta}(s;0.1,0.1)$ as the proposal distribution $\tilde{q}$. As shown in Figure 6, the values of $\frac{1}{\tilde{q}(s_{i})}l(s_{i})$ are plotted against their $s$, which becomes more concentrated in a small range.
Figure 6: Comparison between uniform sampling and importance sampling. Each dot represent the loss of one sample with respect to the diffusion time.
C.4 Additional forecasting results
Exchange rate Solar Traffic Wiki
DDPM 0.011$\pm$0.004 0.377$\pm$0.061 0.064$\pm$0.014 0.093$\pm$0.023
FM 0.011$\pm$0.001 0.445$\pm$0.031 0.041$\pm$0.002 80.624$\pm$89.804
SGM 0.01$\pm$0.002 0.388$\pm$0.026 0.08$\pm$0.053 0.122$\pm$0.026
SI 0.008$\pm$0.002 0.399$\pm$0.065 0.089$\pm$0.006 0.091$\pm$0.011
Table 4: ND-sum. A smaller number indicates better performance.
Exchange rate Solar Traffic Wiki
DDPM 0.013$\pm$0.005 0.72$\pm$0.08 0.094$\pm$0.029 0.123$\pm$0.026
FM 0.014$\pm$0.002 0.849$\pm$0.072 0.059$\pm$0.007 165.128$\pm$147.682
SGM 0.019$\pm$0.004 0.76$\pm$0.066 0.109$\pm$0.064 0.164$\pm$0.03
SI 0.01$\pm$0.003 0.722$\pm$0.132 0.127$\pm$0.003 0.117$\pm$0.011
Table 5: NRMSE-sum. A smaller number indicates better performance.
C.5 Baseline Model using Unconditional SI
Additionally, we introduced a new experiment to verify the necessity of conditional SI over unconditional SI. The unconditional SI diffuses from pure noise and does not utilize the prior distribution from the previous time point. In this case, the context for the prediction is provided exclusively by the RNN encoder. The new results are shown in the following tables. When compared with the conditional SI framework, the unconditional model shows slightly inferior performance.
Exchange rate Solar Traffic
SI 0.007$\pm$0.001 0.359$\pm$0.06 0.083$\pm$0.005
Vanilla SI 0.010$\pm$0.001 0.383$\pm$0.010 0.082$\pm$0.006
Table 6: CRPS-sum metric on multivariate probabilistic forecasting. A smaller number indicates better performance.
Exchange rate Solar Traffic
SI 0.008$\pm$0.002 0.399$\pm$0.065 0.089$\pm$0.006
Vanilla SI 0.010$\pm$0.003 0.430$\pm$0.113 0.093$\pm$0.007
Table 7: ND-sum. A smaller number indicates better performance.
Exchange rate Solar Traffic
SI 0.010$\pm$0.003 0.722$\pm$0.132 0.127$\pm$0.003
Vanilla SI 0.012$\pm$0.003 0.815$\pm$0.135 0.132$\pm$0.015
Table 8: NRMSE-sum. A smaller number indicates better performance.